# Cayley Graph Propagation

### v1

@@ Abstract
Graph neural networks (GNNs) have emerged as a cornerstone for processing graph-structured data [1] with significant contributions in various domains and real-world applications [2, 3]. Nearly all GNNs are dependent on propagating information between neighbouring nodes in the graph [4], whereby the message-passing paradigm [5] serves as an architecture for facilitating this information. This paradigm involves the iterative exchange of messages, with nodes aggregating and updating their representations based on received information from neighbours. However, a phenomenon known as over-squashing [6] can limit the effectiveness of this message-passing process. Over-squashing occurs when a large volume of messages are aggregated into fixed-size vectors, which hinders the expressive power of GNNs, especially when dealing with long-range node interactions [7]. Consequently, the input graph’s topology is a key contributing factor to the over-squashing problem.
@@
For this reason, the over-squashing phenomenon is an active area of research with several techniques proposed to mitigate over-squashing. Several recent works have analysed it through varying lenses, including curvature [8], spectral expansion properties [9, 10] and effective resistance [11, 12]. The majority of solutions to this problem fall under the category of graph rewiring, in which the graph topology is directly modified based on an optimisation target. However, this imposes the computational complexity of having to surgically analyse the graph structure. Notably, Expander Graph Propagation (EGP) [13] propose a unique solution of constructing an independent desirable graph structure to propagate information over.
@@
Moreover, in the EGP paper the authors identified desirable criteria to mitigate over-squashing and effectively handle global context in graph representation learning: global information propagation (i), no bottlenecks (ii), subquadratic time and space complexity (iii) and no dedicated preprocessing (iv). The authors surveyed prior approaches, including traditional GNNs (iii, iv), master-node methods (i, iii, iv) [5, 14], fully connected graphs (i, ii, iv) [6] and the aforementioned graph rewiring techniques (i-iii), ultimately recognising the efficacy of expander graphs [15, 16] as the desirable graph structure for bottleneck-free information propagation.
@@
In the EGP paper, a family of expander graphs have been constructed leveraging the well-known theoretical results of special linear groups, for which a family of corresponding Cayley graphs can be derived. This family of Cayley graphs is guaranteed to have the desirable expansion properties and are precomutable (i-iv). Importantly, although Cayley graphs are scalable, achieving a specific number of nodes is not always feasible; for instance, the number of nodes for a group of $n$ is known to be $O(|V|^{3})$. This consideration is addressed within the paper by identifying the smallest $n$ that yields a graph larger or equal to the desired number of nodes, and then subsequently truncating the Cayley graph to match the input graph’s number of nodes in breadth-first order.
@@
Motivated by the promising research direction of Deac et al. [13] in which the authors use an independent graph structure that is theoretically known to exhibit desirable properties we propose a different approach of using the Cayley graph. We propose a more optimal approach of embracing the complete Cayley graph structure, guaranteeing the coveted expansion properties.
@@
In this paper, we present Cayley Graph Propagation (CGP) [^1] Our source code is available at:https://github.com/josephjwilson/cayley_graph_propagation, a novel model that uses the complete Cayley graph structure, which is still decoupled from the input graph’s topology, fulfilling the desirable criteria as set by EGP (i-iv). Our contributions are as follows:
  • In Section 3, we highlight the theoretical benefits of using Cayley graphs for message propagation. We show that the truncation procedure performed in EGP to align the Cayley graph with the input graph may be detrimental to the coveted theoretical benefits.
@@ 1 Background
Given an undirected graph denoted as $G=(V,E)$, where $V$ and $E$ denote its nodes and edges respectively. The topology of the graph is encoded in the adjacency matrix ${\bf A}\in\mathbb{R}^{|n|\times|n|}$, where the number of nodes $n=|V|$. Let ${\bf D}={\bf D}(G)$ denote the diagonal matrix of degrees as given by ${\bf D}_{vv}=d_{v}$. The normalised graph Laplacian ${\bf L}={\bf L}(G)$ is defined by ${\bf L}={\bf D}^{-1/2}({\bf D}-{\bf A}){\bf D}^{-1/2}$. From the normalised Laplacian $L$ the eigenvalues $0=\lambda_{0}\leq\lambda_{1}\leq...\leq\lambda_{n-2}\leq\lambda_{n-1}$. Importantly, from this derivation the spectral gap of graph $G$ is $\lambda_{1}-\lambda_{0}=\lambda_{1}$; the Cheeger inequality then defines that a larger spectral gap of graph $G$ is an indicator of good spectral expansion properties. Accordingly, a graph with desirable expansion properties (or a larger spectral gap) defines that it has strong connectivity, or alternatively it is globally lacking bottlenecks [13].
@@
Due to the definition of an expander graph, there are consequently several known construction approaches [17, 18]. We will focus on the deterministic algebraic approach as introduced in EGP [13]. A family of expander graphs have been precomputed leveraging the well-known theoretical results of special linear groups, $\mathrm{SL}(2,\mathbb{Z}_{n})$, for which a family of corresponding Cayley graphs, $\mathrm{Cay}(\mathrm{SL}(2,\mathbb{Z}_{n});S_{n})$, can be derived. Here, $S_{n}$ ([13], Definition 8) denotes a particular generating set for $\mathrm{SL}(2,\mathbb{Z}_{n})$. For appropriate choices of $S_{n}$, the corresponding Cayley graphs are guaranteed to have expansion properties. Moreover, from Figure 1 we see that the constructed graph are 4-regular, $(|E|=2|V|)$. Importantly, although Cayley graphs are scalable, achieving a specific number of nodes is not always feasible; for instance, the node count of $\mathrm{Cay}(\mathrm{SL}(2,\mathbb{Z}_{n});S_{n})$ is given as per:
@@
Over-smoothing
Independent from over-squashing, but another well-known problem impacting the expressivity of GNNs is over-smoothing [21, 22]. This phenomenon occurs in GNNs when the number of layers increases [23, 24], such that node features become increasingly similar [25]. However, the over-smoothing problem is correlated with over-squashing due to a common approach to the latter being graph rewiring; too many additional edges lead to over-smoothing [9]. There are varying approaches to measure over-smoothing for a graph with one such notable metric being the Dirichlet energy [26, 9, 12].
@@ 2 Existing approaches to mitigate over-squashing
In this part, we explore the current landscape of several novel techniques that try to alleviate the over-squashing phenomenon [6]. In essence, the main principle behind many of these techniques is to decouple the input graph $G$ from the computational one, such that it has structurally fewer bottlenecks. Alon and Yahav [6] simply proposed a rewiring technique that does not require the analysis of the input graph by making the last layer of the GNN fully adjacent (FA), allowing all nodes to interact with each other. The effectiveness of such an approach can be shown by Graph Transformers [27, 28], where every layer is fully-connected. However, such an approach is limited by even modest graph sizes due to it imposing $O(|V|^{2})$ edges. An alternative approach is a master node [5]; here a new node is introduced, which is connected to all of the nodes within the graph. This approach is effective as it reduces the graph’s diameter to $2$ by only adding one new node with $O(|V|)$ edges. However, the master node itself becomes the bottleneck. Notably, both of the aforementioned approaches are independent in relation to the input graph topology, therefore satisfying (iv) of no dedicated preprocessing. Conversely, an approach towards this phenomenon is graph rewiring techniques [9, 12, 8], in which the input graph’s connectivity is altered in accordance with an optimisation goal.
@@
Even though Alon and Yahav [6] were the first to introduce the over-squashing phenomenon, a plethora of research has formalised this problem [8, 10, 20]. To this end, an abundance of graph rewiring techniques have stemmed to modify the graph connectivity to try and mitigate bottlenecks. A popular class of approaches are based on a spectral quantity of a graph [9] or to reduce the effective resistance [12, 10, 11]. These approaches have provided promising insights and have empirically reduced over-squashing, but they impose a computational complexity of having to examine the input graph structure.
@@ 3 Benefits of Cayley graphs
Figure 2: Illustrates the correlation between the Cheeger constant (left) and diameter (right) from truncated Cayley graphs of $\mathrm{SL}(2,\mathbb{Z}_{n})$. The dotted red-line represent complete Cayley graph structure for the groups of $\mathbb{Z}_{n}.$
In this section, we further explore the benefits of using Cayley graphs and why they have been used as a conduit in a number of over-squashing approaches [10, 29, 13]. In particular, we explore the implications of the expansion properties in regards to the truncated Cayley graphs used by Deac et al. [13]. In contrast to this, we examine the benefits of using the complete and regular Cayley graph structure through the lens of overfitting as per the recent work of Bechler-Speicher et al. [32].
3.1 Cayley graphs expansion properties
The structure of a Cayley graph makes it an ideal conduit for the propagation of information due to its sparsity and being highly connected. In particular, the family of Cayley graphs derived from $\mathrm{Cay}(\mathrm{SL}(2,\mathbb{Z}_{n});S_{n})$ are in the magnitude of $O(|V|^{3})$, and are well-known to have a high spectral gap. As from Section 1, a higher spectral gap is a desirable expansion property, meaning that it is highly connected. Correspondingly, we conjecture that truncating a Cayley graph to match an input graph as done by Deac et al. [13] will have an adverse impact on the desired expansion properties. In Figure 2 we empirically measure this impact using the Cheeger constant, as well as in respect to the diameter. In Appendix (Section D) we extend this analysis through the perspective of effective resistance, which provides a different approach to quantify over-squashing.
In the pursuit of this, the recent work of Sterner et al. [33] proposes to heuristically align the Cayley graph, such that it better aligns the corresponding edges. However, this distils to a NP-hard problem, so they opt to implement a greedy strategy to align both graphs. This method imposes a costly preprocessing time which does not comply with our desiderata (i-iv). Consequently, we do not pursue this approach further, but this variant could present an interesting avenue for future work.
@@ 3.1 Cayley graphs expansion properties
The diameter of a graph influences the effectiveness of traversal between nodes. A lower diameter facilitates a more efficient graph structure, enabling nodes to reach each other in a shorter number of hops. Our constructed Cayley graph with $|V|$ nodes has a low diameter, requiring only $O(log(|V(G_{i})|)$ ([13], Theorem 5) steps to globally propagate information. This coincides with the work of Di Giovanni et al. [20], as they prove that over-squashing occurs between nodes with a high commute-time. Due to expander graphs having a low diameter, any node within the graph can reach another node in a short number of hops. The results in Figure 2 in relation to the diameter correlate with those shown for the Cheeger constant; the lower diameter is at each complete Cayley graph structure.
3.2 Cayley Graphs as Regular Graphs
In this subsection, we highlight additional benefits of the Cayley graph.
Robustness to Graph Overfitting
It was observed in [32] that graph neural networks tend to overfit the given graph structure, even in cases where it does no provide useful information for the predictive task. Nonetheless, it was shown that regular graphs exhibit robustness to this overfitting. As Cayley graphs are regular graphs, they exhibit this robustness.
We used four different datasets from this baseline by sampling graph-structures from different graph distributions. The set of node feature vectors remains the same across all the datasets, and thus, the datasets differ only in their graph structure. The graph distributions we used are: Cayley graphs over $24$ nodes, star-graph (Star) where the only connections are between one specific node and all other nodes, and the preferential attachment model (BA)  [37], where the graph is built by incrementally adding new nodes and connecting them to existing nodes with probability proportional to the degrees of the existing nodes. We used the data as is with empty graphs (Empty) as a baseline to compare to.
On each dataset, we varied the training set size and evaluated test errors on $5$ runs with random seeds. More details on the experiment can be found in the Appendix (Section F).
Extrapolation
The ability and failures of GNNs to extrapolate to graphs of sizes larger then the one presented during training was examined in Yehudai et al. [38]. It was shown that size generalisation is dependent on local structures around each node, called d-patterns. In particular, if increasing the graph size does not change the distribution of these d-patterns, then extrapolation to larger graph sizes is guaranteed. In the context of Cayley graphs, this implies the following lemma.
Lemma 3.1.
Let $D_{G}$ be a distribution over $\mathrm{Cay}(\mathrm{SL}(k,\mathbb{Z}_{n});S_{n})$ and $D_{X}$ be a distribution over node features. Assume a node classification task with training set of infinite size sampled from $D_{G}$ and $D_{X}$. Denote the learned model by $f$. Assume that test examples are sampled from $D^{\prime}_{G}$, a distribution over $\mathrm{Cay}(\mathrm{SL}(k,\mathbb{Z}_{n^{\prime}});S_{n^{\prime}})$ where $n\neq n^{\prime}$ and $D_{X}$. Then $f$ will have zero test error.
The proof is provided in the Appendix (Section G) and it utilises the property that increasing the modulus $n$, increases the Cayley graph while preserving the regularity degree.
@@ 4 Cayley Graph Propagation
In the previous sections, we provided theoretical motivations for our proposed method of utilising the complete Cayley graph structure. The setup for CGP closely aligns with that of EGP in most aspects. We continue to consider the input to a GNN as a node feature matrix ${\bf X}\in\mathbb{R}^{|V|\times d}$ and an adjacency matrix ${\bf A}\in\mathbb{R}^{|V|\times|V|}$, which can be fed in an edge-list manner.
Additionally, the construction of the Cayley graph $\mathrm{Cay}(\mathrm{SL}(2,\mathbb{Z}_{n});S_{n})$ is still done by choosing the smallest $n$ such that $|V(\mathrm{Cay}(\mathrm{SL}(2,\mathbb{Z}_{n});S_{n}))|\geq|V|$. However, we no longer truncate the Cayley graph such that a subgraph ${\bf A}^{\text{Cay}(n)}_{1:|V|,1:|V|}$ is extracted – instead, we opt for a different approach of retaining all of the nodes of the Cayley graph, and its corresponding adjacency matrix ${\bf A}^{\text{Cay}(n)}$.
@@
where $\mathrm{InitVirt}$ is any sampling procedure for initialising $d$-dimensional feature vectors; for example, we may choose to sample random features from $\mathcal{N}(0,1)$, or in our case initialise them to zeros.

### v2

@@ Abstract
Graph neural networks (GNNs) have emerged as a cornerstone for processing graph-structured data [1] with significant contributions in various domains and real-world applications [2, 3]. The majority of GNNs are dependent on propagating information between neighbouring nodes in the graph [4], known as Message Passing Neural Networks (MPNNs) [5]. This message-passing paradigm involves the iterative exchange of messages, with nodes aggregating and updating their representations based on received information from their neighbours. Thus, a sufficient number of layers are required to be able to capture long-range interactions. However, increasing the number of layers results in an exponential growth of the receptive field and subsequently leading to a large volume of messages to be aggregated into fixed-sized vectors. This phenomenon is known as over-squashing [6] and hinders the expressivity of GNNs [7]. The underlying graph topology has been proven to be a key contributing factor to the over-squashing problem [8].
@@
The over-squashing phenomenon is an active area of research with several techniques proposed to alter the topological properties of an input graph to mitigate over-squashing. Several recent works have analysed it through varying lenses, including curvature [9], spectral expansion properties [10, 11] and effective resistance [12, 13]. Most of the solutions to this problem fall into the category of graph rewiring, in which the graph topology is directly modified based on an optimisation target. However, this imposes the computational complexity of having to surgically analyse the graph structure. Notably, the recent work of Expander Graph Propagation (EGP) by Deac et al. [14] propose a unique solution of constructing an independent desirable graph structure to propagate information over.
@@
To this end, Deac et al. [14] identified four desirable criteria to mitigate over-squashing and effectively handle global context in graph representation learning: global information propagation (i), no bottlenecks (ii), subquadratic time and space complexity (iii) and no dedicated preprocessing (iv). The authors surveyed prior approaches, including traditional GNNs (iii, iv), master-node methods (i, iii, iv) [5, 15], fully connected graphs (i, ii, iv) [6] and the aforementioned graph rewiring techniques (i-iii). Ultimately, Deac et al. [14] recognised the efficacy of expander graphs [16, 17] as the desirable graph structure for bottleneck-free information propagation, due to their favourable topological properties.
@@
A family of expander graphs in Deac et al. [14] has been constructed leveraging the well-known theoretical results of special linear groups, for which a family of corresponding Cayley graphs can be derived. This family of Cayley graphs is guaranteed to have the derisible topological properties to mitigate over-squashing, as well as being efficiently precomputable (i-iv). Although the constructed Cayley graphs are scalable, the number of nodes are in order of $O(|V|^{3})$. This consideration is addressed within EGP [14] by identifying the smallest $n$ that yields a graph larger or equal to the desired number of nodes, and then subsequently truncating the Cayley graph to match the input graph’s number of nodes in breadth-first order. Consequently, this truncation procedure results in a subgraph derived from the expander graph being used as the computational template.
@@
Motivated by the promising research direction of Deac et al. [14] in which the authors use an independent graph structure that is theoretically known to exhibit desirable properties, we introduce an alternative approach of using the Cayley graph. We propose a more optimal approach of embracing the complete Cayley graph structure, guaranteeing the coveted topological properties.
@@
In this paper, we present Cayley Graph Propagation (CGP) [^1] Our source code is available at:https://github.com/josephjwilson/cayley_graph_propagation, a novel model that uses the complete Cayley graph structure to mitigate over-squashing, whilst still fulfilling the derisible criteria as set by Deac et al. [14] (i-iv). Our contributions are as follows:
  • In Section 3, we highlight the optimal topological properties of Cayley graphs for message propagation. We show that the truncation procedure performed in EGP to align the Cayley graph with the input graph may be detrimental to the coveted theoretical benefits.
@@ 1 Background
Given an undirected graph denoted as $G=(V,E)$, where $V$ and $E$ denote its nodes and edges respectively. The topology of the graph is encoded in the adjacency matrix ${\bf A}\in\mathbb{R}^{|V|\times|V|}$, where $|V|$ is the number of nodes. Let ${\bf D}={\bf D}(G)$ denote the diagonal matrix of degrees as given by ${\bf D}_{vv}=d_{v}$. The normalised graph Laplacian ${\bf L}={\bf L}(G)$ is defined by ${\bf L}={\bf D}^{-1/2}({\bf D}-{\bf A}){\bf D}^{-1/2}$. From the normalised Laplacian ${\bf L}$ the eigenvalues $0=\lambda_{0}\leq\lambda_{1}\leq...\leq\lambda_{n-2}\leq\lambda_{n-1}$. Importantly, from this derivation the spectral gap of graph $G$ is $\lambda_{1}-\lambda_{0}=\lambda_{1}$; the Cheeger inequality then defines that a larger spectral gap of graph $G$ is an indicator of good spectral expansion properties. Accordingly, a graph with desirable expansion properties (or a larger spectral gap) defines that it has strong connectivity, or alternatively it is globally lacking bottlenecks [14].
@@
Due to the definition of an expander graph, there are consequently several known construction approaches [18, 19]. We will focus on the deterministic algebraic approach as introduced by Deac et al. [14]. A family of expander graphs have been precomputed leveraging the well-known theoretical results of special linear groups, $\mathrm{SL}(2,\mathbb{Z}_{n})$, for which a family of corresponding Cayley graphs, $\mathrm{Cay}(\mathrm{SL}(2,\mathbb{Z}_{n});S_{n})$, can be derived. Here, $S_{n}$ ([14], Definition 8) denotes a particular generating set for $\mathrm{SL}(2,\mathbb{Z}_{n})$. For appropriate choices of $S_{n}$, the corresponding Cayley graphs are guaranteed to have expansion properties. Moreover, from Figure 1 we see that the constructed graph are 4-regular, $(|E|=2|V|)$. Importantly, although Cayley graphs are scalable, achieving a specific number of nodes is not always feasible; for instance, the node count of $\mathrm{Cay}(\mathrm{SL}(2,\mathbb{Z}_{n});S_{n})$ is given as per:
@@ 2 Existing approaches to mitigate over-squashing
In this section, we explore the current landscape of several novel techniques that try to alleviate the over-squashing phenomenon [6]. In essence, the main principle behind many of these techniques is to decouple the input graph $G$ from the computational one, such that it has structurally fewer bottlenecks. Alon and Yahav [6] simply proposed a rewiring technique that does not require the analysis of the input graph by making the last layer of the GNN fully adjacent (FA), allowing all nodes to interact with each other. The effectiveness of such an approach can be shown by (dense) Graph Transformers [21, 22], where every layer is fully-connected. However, such an approach is limited by even modest graph sizes due to it imposing $O(|V|^{2})$ edges. An alternative approach is a master node [5]; here a new node is introduced, which is connected to all of the nodes within the graph. This approach is effective as it reduces the graph’s diameter to $2$ by only adding one new node with $O(|V|)$ edges. However, the master node itself becomes the bottleneck. Notably, both of the aforementioned approaches are independent in relation to the input graph topology, therefore satisfying (iv) of no dedicated preprocessing.
@@
An alternative promising approach line of research is to rewire the input graph $G$ to optimise the spectral or spatial properties of the graph. To this end, an abundance of graph rewiring techniques have stemmed to modify the graph connectivity to try and mitigate bottlenecks. A popular class of approaches are based on a spectral quantity of a graph [10] or to reduce the effective resistance [13, 11, 12]. These approaches have provided promising insights and have empirically reduced over-squashing, but they impose a computational complexity of having to examine the input graph structure.
@@ 3 Benefits of Cayley graphs
Figure 2: Illustrates the correlation between the Cheeger constant (left) and diameter (right) between extracting a truncated Cayley graphs from $\mathrm{SL}(2,\mathbb{Z}_{n})$. The properties exhibit the most derisible property at each given complete Cayley graph interval (as denoted by the dotted red line) for the groups of $\mathbb{Z}_{n}$. The trend shows a detrimental drop in the derisible property occurring once the truncated Cayley graph aligns with the next complete Cayley graph structure.
In this section, we formalise the topological properties of expander graphs and why they have been used as a conduit in a number of over-squashing approaches [11, 23, 14]. In particular, we focus on the Cayley graph expander family as constructed by Deac et al. [14]. Furthermore, we importantly anaylse the impact on the topological properties of the truncated Cayley graphs as used by Deac et al. [14]. In contrast to this, in the appendix, we examine an alternative benefit to Cayley graphs being regular graphs through the lens of the recent work of Bechler-Speicher et al. [26].
3.1 Topological properties of Cayley graphs
The Cayley graph’s topology serves as ideal conduit for the propagation of information due to its sparsity and being highly connected. In particular, we recall that the family of Cayley graphs derived from $\mathrm{Cay}(\mathrm{SL}(2,\mathbb{Z}_{n});S_{n})$ are in the magnitude of $O(|V|^{3})$, and are well-known to exhibit a high spectral gap, low diameter and optimal commute time.
@@ 3.1 Topological properties of Cayley graphs
The diameter of a graph influences the effectiveness of traversal between nodes. Intuitively, a lower diameter facilitates a more efficient graph structure, enabling nodes to reach each other in a shorter number of hops. Our constructed Cayley graph with $|V|$ nodes has a low diameter, requiring only $O(log(|V(G_{i})|)$ ([14], Theorem 5) steps to globally propagate information. The results in Figure 2 in relation to the diameter correlate with those shown for the Cheeger constant; the lower diameter is at each complete Cayley graph structure.
Commute time
The recent work of Di Giovanni et al. [7, 8] validates that the topology of the graph is the primary factor affecting over-squashing. Theorem 5.5 from Di Giovanni et al. [7] establishes that the extent of over-squashing between a pair of nodes $u$ and $v$ over-squashing can be bounded by the commute time $T(u,v)$. The commute time is defined as the expected number of steps in a random walk from $u$ to $v$ and return back to $u$. Consequently, over-squashing occurs between nodes with a large commute time. Notably, the family of Cayley graphs as derived by Deac et al. [14] are directly mentioned by Di Giovanni et al. [7, 8] as being the optimal computational template for message-passing due to the commute time scaling linearly with the number of edges. We note that the commute-time $T(u,v)$ between a pair of nodes is analogous to the effective resistance [12]. In Appendix (Section A), we extend our analysis of the impact of truncating a Cayley graph to align with an input graph through the perspective of effective resistance.
@@ 4 Cayley Graph Propagation
In the previous sections, we provided theoretical motivations for our proposed method of utilising the complete Cayley graph structure. The setup for CGP closely aligns with that of EGP in most aspects. We consider the input to a GNN as a node feature matrix ${\bf X}\in\mathbb{R}^{|V|\times d}$ and an adjacency matrix ${\bf A}\in\mathbb{R}^{|V|\times|V|}$.
The construction of the Cayley graph $\mathrm{Cay}(\mathrm{SL}(2,\mathbb{Z}_{n});S_{n})$ is done by choosing the smallest $n$ such that $|V(\mathrm{Cay}(\mathrm{SL}(2,\mathbb{Z}_{n});S_{n}))|\geq|V|$. However, we no longer truncate the Cayley graph such that a subgraph ${\bf A}^{\text{Cay}(n)}_{1:|V|,1:|V|}$ is extracted – instead, we opt for a different approach of retaining all of the nodes of the Cayley graph, and its corresponding adjacency matrix ${\bf A}^{\text{Cay}(n)}$.
@@
where $\mathrm{InitVirt}$ is any sampling procedure for initialising $d$-dimensional feature vectors. For example, we may choose to sample random features from $\mathcal{N}(0,1)$, or in our case initialise them to zeros [5, 30, 31]. In Appendix (Section B), we provide additional studies to compare the performance of different initialisation strategies.